OpenStudy (anonymous):

What is the exluded values of the following functions? 1. y = 4/(9x - 45) 2. y = 7/(2x + 48)

OpenStudy (anonymous):

excluded values are like holes on a function or graph. which tend to happen when the denominator equals to 0

Parth (parthkohli):

You get the excluded value when you get 0 as the denominator. First one- \(\Large \color{MidnightBlue}{\Rightarrow 9x - 45 = 0 }\) Solve for x to get the excluded value. Same for second.

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow 2x + 48 = 0 }\) Here's for the second

OpenStudy (anonymous):

The value of denominator should not be equal to zero as the value of function would become infinity. Therefore 9x−45=0 -->The value that x should not be equal to

OpenStudy (anonymous):

0 is not an answer choice.

Parth (parthkohli):

Solve for x in the equations I have given.

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow 9x = 45 }\) \(\Large \color{MidnightBlue}{\Rightarrow x = 5 }\) Therefor, excluded value in the first one is 5.

OpenStudy (anonymous):

right, but what we mean is that to find the excluded values you need to set the denominator to 0. for example the first problem just taking the denominator into consideration since its what we are looking at 9x+45=0 9x=-45 x=-45/9

OpenStudy (anonymous):

*** sorry, @ParthKohli is right, I had my number wrong should be 9x-45 which will change the sign to positive. sorry about that